OpenStudy (anonymous):

what is the limit for x=(5+x)^0.5

Parth (parthkohli):

x approaches what?

OpenStudy (anonymous):

infinit

Parth (parthkohli):

\[\large \rm (5 + x)^{0.5} = \sqrt{5 + x}\]

Parth (parthkohli):

It's just infinite.

Parth (parthkohli):

No manipulation can be done in this case.

Parth (parthkohli):

Because\[\large \rm \lim_{x\to \infty} \sqrt{x + 5} = \sqrt{\lim_{x \to \infty} (x + 5)}\]

OpenStudy (anonymous):

maybe is Xn+1=(Xn + 1 )^0.5

OpenStudy (anonymous):

its a sequence ( I already proved its bounded and monotone